# Automated Retinal Layer and Fluid Segmentation and Cross-sectional Analysis using Spectral Domain Optical Coherence Tomography Images for Diabetic Retinopathy

### v1

@@ III-B Thickness
Figure 7: The distribution of layer thickness and fluid volume measurements across different sectors for patients diagnosed with NPDR and PDR. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles. The diamond markers represent the mean thickness for each group, calculated from the original data without outlier removal.
(a) Example of 54-year-old male patients with right eye image.
(b) Example of 57-year-old male patients with left eye image.
@@
Figure 9: Examples of ETDRS thickness map comparison between NPDR and PDR patient with matched age, gender and eyeside. For each region, the first row shows the thickness heatmap overlayed onto the layer En Face image. The second row shows the quantitative average thickness for each sector.
We also evaluated the association between retinal layer sector thickness and visual acuity, as shown in table IV. The evaluation is performed for NPDR and PDR separately. Each value represents the Pearson correlation coefficient with a p-value within a specific layer sector. The p-values are adjusted by age, gender and diabetes duration. For NPDR, significant negative correlations are found in the SI sector of INL plus the II and SO sectors of ONL+IS. It suggests that layer thickening in these regions is associated with better vision. However, in PDR, we observe significant positive correlations in the II, SO, NO and IO sectors of RNFL, SO sector of OPL, NO sector of ONL+IS, TO sector of EZ, multiple sectors for both fluid (II, TI, SO, NO, TO) and HRF (SI, SO, NO). This finding indicates that thickening these layers and accumulations of fluid and HRF may result in worse vision.
Discussion
@@
The segmentation performance varied across models, highlighting differences in architectural strengths and their ability to segment specific retinal layers and fluid-related abnormalities. SwinUNETR and VM-UNet consistently achieved high DSC and NSD scores, indicating their robustness in handling complex retinal structures. SwinUNETR particularly excelled in segmenting the OPL and HRF layers, which may be attributed to its transformer-based architecture that effectively captures long-range dependencies. VM-UNet, on the other hand, performed better in segmenting the fluid regions, suggesting that its sequential nature enhances segmentation continuity, particularly in areas with less distinct boundaries.
@@
The performance differences in DSC and NSD indicate that while both models perform well, their strengths lie in different layers. DSC reflects volumetric segmentation accuracy, whereas SwinUNETR demonstrated better performance in HRF, likely due to its ability to differentiate small scattered hyper-reflective lesions. However, NSD provides insights into surface delineation accuracy, where VM-UNet showed superior performance in fluid regions, maintaining smoother boundaries compared to other models. U-Net and SegFormer, though competitive in some layers, exhibited weaker performance in fine layer segmentation, particularly in RNFL and OPL, where structural continuity is essential for accurate disease characterization. Although the SwinUNETR slightly outperforms VM-UNet in several regions, VM-UNet has significantly lower computational complexity (O(N)) than SwinUNETR (O($N^{2}$), which is crucial for remote deployment in clinics with limited computational resources.
@@
The segmentation of fluid and HRF remains a significant challenge across all models. Fluid regions exhibit substantial variability, with VM-UNet demonstrating better spatial continuity but often under-segmenting these regions. In contrast, SwinUNETR captures fluid regions more extensively but is prone to occasional over-segmentation. HRF segmentation presents an even more significant challenge due to the presence of small, widely distributed hyper-reflective regions. Both models tend to under-segment fluid and HRF, frequently misclassifying them into adjacent retinal layers such as OPL and ONL+IS. Moreover, SwinUNETR generally exhibits a greater tendency to under-segment retinal regions than VM-UNet. Although significant weight adjustments were applied to fluid and HRF regions, as described in Section II-E, additional strategies are needed to enhance model learning and improve segmentation performance in these complex regions.
Significant differences in retinal layer thickness between NPDR and PDR offer valuable insights into the progression of DR. Prior to adjusting for potential confounders, including age, gender, and duration of diabetes, RNFL was significantly thicker in the NI and TI sectors in PDR. This finding contrasts with previous reports indicating RNFL thinning in more advanced stages of NPDR and PDR (Ulhaq et al., 2022)(Bhaskaran et al., 2023). It suggests that transient swelling, likely induced by abnormal blood vessel growth, may occur during the early acute phase of PDR but subsequently diminishes as chronic edema and neuroinflammation lead to tissue loss (Garcia et al., 2013). After adjusting for confounders, statistical significance persists in these two RNFL sectors, suggesting that these structural changes alone may serve as potential biomarkers for the early detection of PDR. When comparing these findings to model predictions, VM-UNet demonstrates greater consistency with ground-truth results than SwinUNETR, particularly in the RNFL and OS sectors. The loss of statistical significance for HRF is likely due to under-segmentation by both models. This highlights the sensitivity of retinal thickness analysis to segmentation variations and underscores VM-UNet’s suitability for detecting significant differences in retinal thickness between NPDR and PDR.
GCL+IPL exhibited significant thinning in PDR in the II, TI, and IO sectors before adjustment, reflecting the apoptosis of ganglion cells due to potential neurodegeneration and inflammatory responses. However, after adjusting for confounders, the TI and IO sectors are no longer significant, suggesting that systemic factors play a role in some but not all aspects of GCL+IPL thinning in PDR. Similarly, the OS layer showed considerable thinning in PDR across all sectors before confounders adjustment, highlighting widespread photoreceptor degeneration and atrophy. After adjustment, thinning in the TI sector lost significance. This suggests that overall OS deterioration is a hallmark of DR progression and an early marker of photoreceptor damage in PDR.
While not statistically significant in the t-test analysis, fluid accumulation was more prevalent in PDR than NPDR based on example cases shown in Figure 9. This suggests that while variability in fluid distribution may limit statistical significance, extensive fluid in PDR remains clinically relevant. The increased fluid volume aligns with the pathological features of PDR, where extensive vascular leakages lead to blood-retina barrier breakdown and, eventually, fluid accumulation. Before adjustment, HRF was significantly more widespread in NPDR than PDR, particularly in the II and SO sectors, supporting the hypothesis that HRF is an early marker of inflammation. However, this does not necessarily signify reduced disease severity of PDR but rather a transition to neovascularization toward deeper retinal layers. The chronic vascular leakage may cause the HRF to disperse or form into larger hard exudates and hemorrhages with less visibility in OCT. After adjusting for confounders, these differences lost statistical significance. This suggests that while HRF may be an important early indicator of neuroinflammation, its distribution may be influenced by demographic and disease-specific factors.
Our findings reveal significant associations between retinal layer thickness and visual acuity within NPDR and PDR patients. In NPDR, the observed negative correlations between thickness in specific retinal sectors and visual acuity suggest that moderate thickening in specific layers may have a protective effect with latencies. Conversely, positive correlations across multiple layers in PDR indicate that further increased thickness without proper treatment will lead to vision decline. Specifically, in NPDR, negative correlations in the SI sector of the INL and II/SO sectors of the ONL+IS suggest that mild thickening in these regions may represent a compensatory or adaptive response to metabolic activity, potentially maintaining structural integrity and delaying visual deterioration.
For PDR, thickening in several RNFL sectors (II, SO, NO, IO) is associated with poorer vision. This aligns with previous findings that transient edema and axonal swelling may occur in the early stages of PDR, coinciding with significant visual impairment. Additionally, increased thickness in the SO sector of the OPL, NO sector of the ONL+IS, and TO sector of the EZ is strongly linked to worse vision, likely due to edema-induced synaptic dysfunction between photoreceptors and bipolar cells, as well as EZ disruption affecting photoreceptor integrity.
Furthermore, fluid and HRF accumulation strongly correlate with PDR vision loss but not NPDR. The increased presence of fluid (II, TI, SO, NO) and HRF (SI, SO, NO) suggests progressive neurodegeneration and inflammatory activity in PDR, reinforcing the need for early intervention during the NPDR stage to prevent irreversible retinal damage.
While these findings provide valuable insights into the structural and functional relationships in NPDR and PDR, they are inherently result-oriented and restricted to the specific cohort analyzed. Future research should incorporate larger, multi-center cohorts with diverse demographic and clinical characteristics to enhance their generalizability. Validations in the interventional clinical trials will also help determine their potential applicability in guiding personalized treatment strategies for DR.
IV Limitation and Future Work
@@
First, manual labelling imperfections impact both model performance and thickness measurements. Retinal layer segmentation is inherently challenging due to subtle boundary variations and overlapping structures. In cases where excessive fluid penetrates the layer boundaries, some portions of the layer become invisible or physically diminished. More clinical expertise is needed to segment the extreme instances properly. HRF segmentation suffers from inconsistencies in ground-truth annotations, as small, widely distributed foci are challenging to delineate manually. Interestingly, in some cases, automated models provided more precise segmentations than the manual ground truth, particularly for HRF regions. Figure 10 shows that the model prediction outperforms the ground truth in HRF segmentation. This suggests that deep learning models could be leveraged for segmentation and refining human annotations. Additionally, using more pre-processing and post-processing techniques may help improve the performance, such as the pixel-wise relative positional map as an extra input and random forest classifier as a label refiner(Ma et al., 2021).
@@
Third, this study’s cross-sectional nature limits its ability to track disease progression over time. Longitudinal studies would provide better insights into how retinal layer thickness evolves in DR. For example, this study observes RNFL thickening in PDR, but longitudinal research often reports RNFL thinning due to neurodegeneration. This suggests that the transient swelling observed in cross-sectional studies may eventually lead to structural atrophy, which can only be confirmed through long-term follow-up studies.
Fourth, while the sample size is sufficient to detect significant differences, it may limit the generalizability of the findings. A larger dataset encompassing a broader range of DR severities and treatment histories could provide more robust conclusions. Expanding the cohort to include more diverse patient populations would also improve the applicability of the findings across different clinical settings.
@@
Fifth, the lack of a detailed NPDR severity grading system may limit the ability to distinguish early, intermediate, and severe NPDR stages. Different NPDR severities likely exhibit distinct retinal layer changes, and a more granular classification system could enhance the understanding of DR progression. Future studies should explore integrating severity-based stratification to assess how thickness variations differ across NPDR subtypes. To our best knowledge, most DR grading datasets with public access focus on fundus color images like Messidor [^3] https://www.adcis.net/en/third-party/messidor/ and DRTiD [^4] https://github.com/FDU-VTS/DRTiD.
@@
Lastly, integrating multi-modal imaging techniques such as OCT angiography (OCTA) could provide additional insights into the vascular changes associated with DR. For example, Alam et al. discovered the difference of vascular complexity features between NPDR and PDR patients(Alam et al., 2021). Multiple OCT parameters are significantly correlated with DR severity(Laotaweerungsawat et al., 2020). Combining structural OCT findings with functional vascular imaging may improve disease characterization and facilitate more targeted therapeutic interventions.
Region Sector
SI NI II TI SO NO IO TO
Non-Proliferative Diabetic Retinopathy (NPDR)
RNFL 0.0093 (0.9424) 0.0329 (0.7983) 0.0857 (0.5042) 0.0270 (0.8338) -0.0083 (0.9488) 0.0532 (0.6791) 0.0223 (0.8623) -0.0325 (0.8005)
GCL+IPL -0.0612 (0.6335) 0.0450 (0.7265) 0.0225 (0.8608) -0.0434 (0.7357) -0.0862 (0.5019) -0.0023 (0.9860) -0.0785 (0.5406) 0.0342 (0.7904)
INL -0.3464 (0.0054) -0.1157 (0.3666) -0.0615 (0.6320) -0.0166 (0.8974) 0.0017 (0.9895) -0.0479 (0.7094) -0.1082 (0.3986) -0.0842 (0.5116)
OPL -0.0642 (0.6171) -0.0515 (0.6883) -0.0777 (0.5452) -0.0341 (0.7905) 0.0825 (0.5204) 0.0779 (0.5440) -0.0854 (0.5056) -0.0636 (0.6206)
ONL+IS -0.1404 (0.2724) -0.0559 (0.6636) -0.2545 (0.0441) -0.0482 (0.7076) -0.2803 (0.0261) -0.0645 (0.6155) -0.1007 (0.4324) -0.0334 (0.7948)
EZ -0.0488 (0.7043) -0.0926 (0.4704) 0.0521 (0.6848) 0.1479 (0.2475) -0.0466 (0.7171) -0.0741 (0.5638) -0.0777 (0.5449) -0.0747 (0.5608)
OS -0.1058 (0.4093) 0.0478 (0.7097) -0.0853 (0.5063) 0.0486 (0.7054) 0.0889 (0.4885) 0.0977 (0.4463) 0.0400 (0.7554) 0.0952 (0.4578)
RPE -0.0816 (0.5249) 0.1665 (0.1922) -0.1138 (0.3743) -0.0633 (0.6219) -0.0143 (0.9117) 0.1426 (0.2650) -0.0028 (0.9826) 0.1983 (0.1192)
Fluid 0.0057 (0.9644) 0.1375 (0.2826) 0.1097 (0.3922) -0.0698 (0.5665) 0.1962 (0.1233) 0.2204(0.0827) 0.0952 (0.4580) -0.0928 (0.4694)
HRF -0.0037 (0.9770) 0.0765 (0.5510) -0.0286 (0.8239) -0.0945 (0.4615) 0.0787 (0.5398) 0.0783(0.5421) -0.0570 (0.6574) -0.1242 (0.3321)
Proliferative Diabetic Retinopathy (PDR)
RNFL 0.1601 (0.2825) 0.1992 (0.1794) 0.2971 (0.0426) 0.2557 (0.0827) 0.3684 (0.0108) 0.3445 (0.0177) 0.3349 (0.0214) 0.1851 (0.2130)
GCL+IPL -0.0184 (0.9024) -0.0757 (0.6129) -0.1019 (0.4953) -0.0155 (0.9177) -0.0330 (0.8259) -0.0793 (0.5963) -0.0604 (0.6867) -0.0456 (0.7607)
INL 0.1897 (0.2017) 0.0554 (0.7113) -0.0196 (0.8958) 0.1538 (0.3020) 0.0125 (0.9333) 0.2325 (0.1157) -0.0116 (0.9384) -0.0001 (0.9995)
OPL 0.1846 (0.2142) 0.2647 (0.0721) 0.1235 (0.4080) 0.2485 (0.0922) 0.3036 (0.0380) 0.2714 (0.0650) 0.2101 (0.1564) 0.2221 (0.1336)
ONL+IS 0.2279 (0.1234) 0.2712 (0.0652) 0.0472 (0.7528) 0.2578 (0.0802) 0.2041 (0.1688) 0.4315 (0.0025) 0.0320 (0.8311) 0.2863 (0.0511)
EZ 0.1297 (0.3850) 0.0653 (0.6628) 0.0527 (0.7250) 0.0458 (0.7596) 0.1073 (0.4730) 0.0588 (0.6946) 0.2042 (0.1686) 0.3334 (0.0220)
OS 0.0194 (0.8968) 0.0340 (0.8204) -0.0568 (0.7047) 0.0318 (0.8321) -0.0492 (0.7428) -0.0093 (0.9505) 0.1254 (0.4011) 0.1112 (0.4566)
RPE -0.0499 (0.7393) -0.0158 (0.9162) 0.0683 (0.6481) -0.0629 (0.6744) -0.0696 (0.6418) 0.1430 (0.3376) 0.2180 (0.1410) 0.2236 (0.1308)
Fluid 0.1768 (0.2346) 0.1607 (0.2806) 0.3300 (0.0235) 0.2877 (0.0499) 0.3010 (0.0398) 0.2914 (0.0469) 0.2255 (0.1275) 0.3372 (0.0204)
HRF 0.3477 (0.0166) 0.2440 (0.0984) 0.1824 (0.2198) 0.2670 (0.0696) 0.3777 (0.0089) 0.3964(0.0058) 0.2549 (0.0837) 0.1004 (0.5020)
TABLE IV: Partial correlation between retinal layer thickness and visual acuity in logMAR for NPDR and PDR. Each cell presents the Pearson correlation coefficient (p-value) adjusted by age, gender, and diabetes duration. Statistically significant values ($p<0.05$) are highlighted in bold.
Figure 10: Example of 62-year-old PDR male patient’s left eye image. The area enclosed by the red rectangle indicates the under-segmentation of HRF from the ground truth, where the model predictions have a better segmentation result.
V Conclusion

### v3

@@ III-B Thickness
Figure 7: The distribution of layer thickness and fluid volume measurements across different sectors for patients diagnosed with NPDR and PDR. The outliers were removed outside the $5^{th}$ and $95^{th}$ percentiles. The diamond markers represent the mean thickness for each group, calculated from the original data without outlier removal.
Figure 8 presents the deviations in retinal thickness measurements from SwinUNETR and VM-UNet segmentations compared to ground truth using GLM. Statistical significance is determined after controlling for multiple comparisons using FDR correction, with filled markers indicating FDR-adjusted p-values below 0.05. SwinUNETR generally demonstrates high agreement with ground truth, with minimal significant deviations except in the RPE layer, particularly in the SI, NI, SO, NO, IO, and TO sectors. In contrast, VM-UNet exhibits more widespread discrepancies, notably in the INL (NI, II, SO, NO, IO, TO), OPL (SI, TI, SO, TO), and EZ (SI, II, SO, NO, IO, TO) layers. Both models yield consistent predictions for fluid and HRF volumes. However, SwinUNETR shows significant overestimation in the SO sector of fluid, and VM-UNet displays significant under-segmentation in the SO and NO sectors of HRF. Additionally, abnormally large CIs are observed in specific sectors, including ONL+IS (TO) and OS (NO) for SwinUNETR, and GCL+IPL (NO) and RPE (TI, TO) for VM-UNet, likely reflecting segmentation failures that lead to extreme thickness estimates.
Figure 8: Coefficient plot showing the statistical difference of thickness measurements between the model predicted segmentation and ground truth based on GLM. The horizontal dashed line represents no difference in retinal thickness. The data with p-value $<$ 0.05 after FDR correction is annotated as filled markers. The data with p-value $>$ 0.05 after FDR correction but not across the reference line is marked as ’*’. Error bars represent 95% confidence intervals.
Figure 9 presents statistical comparisons of retinal layer thickness, fluid volume, and HRF volume between NPDR and PDR groups. The analysis incorporates predicted segmentations from the two top-performing models, SwinUNETR and VM-UNet, alongside ground-truth segmentations for benchmarking. For each layer-sector pair, the GLM was applied to assess the relationship between DR diagnosis and thickness, accounting for potential confounders such as age, gender, and diabetes duration. A positive regression coefficient indicates increased thickness in the PDR group relative to NPDR.
Ground-truth data reveals a significantly increased thickness in the SO sector of the OPL in PDR. Additionally, significant fluid accumulation and HRF presence were observed in the SO and NI sectors, respectively. Conversely, the ONL+IS layer demonstrated significantly greater thickness in the SI and NI sectors in the NPDR group. However, after applying FDR correction, none of these differences remained statistically significant. Both segmentation models exhibited coefficient distributions consistent with the ground truth, although no statistically significant layer-sector differences were observed. Quantitative results are shown in Supplementary Table V
Figure 9: Coefficient plot illustrating the association between NPDR and PDR groups and retinal layer thickness across different retinal sectors, controlling for age, gender, and duration of diabetes. Results were derived from GLM analysis for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents zero effect of DR diagnosis on retinal thickness. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line. No filled markers (statistically significant associations after FDR correction) are present in the graph. Error bars represent 95% confidence intervals.
(a) NPDR
(b) PDR
Figure 10: Coefficient plots illustrating the association between the retinal layer thickness across different retinal sectors and visual acuity (logMAR), controlling for age, gender, and duration of diabetes. The analysis was performed for NPDR shown in 10(a) and PDR shown in 10(b), respectively. Results were derived from generalized linear model regression analyses for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents no effect of retinal thickness on visual acuity. Error bars represent 95% confidence intervals. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line. No filled markers (statistically significant associations after FDR correction) are present in the graph.
Figure 12 presents four examples of thickness comparisons between NPDR and PDR groups using the ETDRS diagram described in Figure 2. Each example shows four regions that are reported with significant thickness differences in Figure 9. Figures 11(a)12(b) correspond to cases of a 57-year-old male OD, a 54-year-old male OS, and a 59-year-old female OD and OS images, respectively. Each pair of patients is matched by age, gender, and eye laterality. For each retinal region, the first row displays the En Face image overlaid with the corresponding layer thickness heatmap, while the second row presents the sector-wise quantitative average thickness. The En Face image is generated using each layer’s maximum intensity projection (MIP). For fluid and HRF, the En Face projection is derived from the entire retinal body (from the ILM to the BM), with thickness representing the accumulated volume in $\mu m^{3}$. From these figures 11(a)12(b), PDR exhibits a larger thickness than NPDR in nearly all OPL sectors. Conversely, the inner sectors of ONL+IS are significantly smaller for PDR. PDR has a larger and broader distribution of fluid and HRF accumulation in most sectors than NPDR. The findings are consistent with previous results.
(a) Example of 57-year-old male patients with right eye image.
(b) Example of 54-year-old male patients with left eye image.
@@
Figure 12: Examples of ETDRS thickness map comparison between NPDR and PDR patients with matched age, gender and eyeside. For each subfigure, only the four significant regions reported in Figure 9 are present. For each region, the NPDR and PDR groups are compared with two representations. The first row shows the thickness heatmap overlayed onto the layer En Face image. The second row shows the quantitative average layer thickness or volume accumulation for each sector.
We further investigated the association between retinal layer-sector thickness and visual acuity (VA), with results summarized in Figure 10. This analysis was conducted separately for NPDR and PDR groups, using the same GLM regression framework applied in Figure 10, adjusting for age, gender, and diabetes duration. Quantitative results are shown in Supplementary Tables VI and VII
In the NPDR group shown in Figure 10(a), ground-truth segmentation revealed that the thickening of several layers was significantly associated with worse visual acuity (higher logMAR values). These included the RNFL (SI), OPL (SO), ONL+IS (SO), and RPE (IO). VM-UNet successfully identified the significant association between OPL thickening in the SO sector and vision loss. Additionally, it reported significant associations in the SO sector of the RNFL and INL, as well as fluid accumulation in the SO sector, all correlating with reduced vision. In contrast, SwinUNETR did not detect any statistically significant associations between retinal layer thickness or pathological volumes (fluid, HRF) and visual impairment. However, after applying FDR correction, none of the associations remained statistically significant.
In the PDR group shown in Figure 10(b), the thickening of several layer sectors was significantly correlated with poorer vision, including the EZ layer (NO) and OS layer (SO, OP and TO). The thinning of the INL (II) was also significantly correlated to worse vision. Both SwinUNETR and VM-UNet correctly identified the significant association between OS thickening and vision loss in the SO and IO sectors and also highlighted a similar association in the NO sector. Furthermore, both models predicted RPE thinning in the II sector, while VM-UNet additionally captured RNFL thinning in the TO sector. However, none of these associations remained statistically significant following FDR correction.
IV Discussion
IV-A Segmentation Models Comparison
@@
This study presents a comprehensive evaluation of the auto-segmentation performance with four state-of-the-art network architectures when segmenting retinal layer, fluid, and HRF on patients that exhibit varying levels of DR severity. The segmentation performance varied across models, highlighting differences in architectural strengths and their ability to segment specific retinal layers and fluid-related abnormalities. Specifically, SwinUNETR and VM-UNet consistently achieved high DSC and NSD scores, indicating their robustness in handling complex retinal structures. SwinUNETR particularly excelled in segmenting the OPL and HRF layers, which may be attributed to its transformer-based architecture that effectively captures long-range dependencies. VM-UNet, on the other hand, performed better in segmenting the fluid regions, suggesting that its sequential nature enhances segmentation continuity, particularly in areas with less distinct boundaries.
@@
The performance differences in DSC and NSD indicate that while both models performed well, their strengths lay in different layers. VM-UNet was superior in several layers plus fluid, whereas SwinUNETR demonstrated better performance in a few layers plus HRF. U-Net and SegFormer, though competitive in some layers, exhibited weaker performance in fine layer segmentation, particularly in RNFL and OPL, where structural continuity is essential for accurate disease characterization. Although the SwinUNETR slightly outperforms VM-UNet in several regions, VM-UNet has significantly lower computational complexity (O(N)) than SwinUNETR (O($N^{2}$)), which is crucial for remote deployment in clinics with limited computational resources.
@@
The segmentation of fluid and HRF remains a significant challenge across all models. Fluid regions exhibit substantial variability, with VM-UNet demonstrating better spatial continuity but often under-segmenting these regions. In contrast, SwinUNETR captures fluid regions more extensively but is prone to occasional over-segmentation. HRF segmentation presents an even more significant challenge due to the presence of small, widely distributed hyper-reflective regions. Both models tend to under-segment fluid and HRF, frequently misclassifying them into adjacent retinal layers such as OPL and ONL+IS. Moreover, SwinUNETR generally exhibits a greater tendency to under-segment retinal regions than VM-UNet. The low DSC and NSD scores for fluid and HRF are mainly caused by their inherently small size relative to the full B-scan and their highly variable shapes and spatial distributions. For most scenarios, these structures occupy only a minor fraction of the retinal cross-section, making their accurate delineation more susceptible to minor boundary deviations. Additionally, their irregular morphology and variable positioning within the retina make consistent segmentation across patients particularly challenging, which disproportionately affects overlap-based metrics despite visually acceptable predictions. Although significant weight adjustments are applied to fluid and HRF regions, as described in Section II-E, additional strategies are needed to enhance model learning and improve segmentation performance in these complex regions.
The accuracy of the predicted segmentation was further evaluated through quantitative analysis of retinal layer thickness. Both SwinUNETR and VM-UNet demonstrated comparable performance, with minimal variation in thickness measurements across most layer-sector combinations when benchmarked against ground-truth annotations. Notably, in this cohort, SwinUNETR outperforms VM-UNet with superior consistency in thickness prediction across most layer sectors except the RPE. These findings suggest that while both models deliver comparable segmentation outputs, subtle segmentation inaccuracies can propagate non-linearly into downstream quantitative metrics such as thickness or volume. Such pixel-level deviations may become magnified in aggregate analyses, potentially leading to misinterpretation in studies relying on precise structural measurements.
Significant differences in retinal layer thickness between NPDR and PDR offer valuable insights into the progression of DR. Ground-truth analysis revealed localized OPL thickening in the SO sector in PDR, likely indicative of extracellular fluid accumulation secondary to microvascular leakage. Similarly, increased fluid volume in the SO sector and HRF burden in the NI sector align with known patterns of retinal inflammation and exudation in advanced DR patients. In contrast, the ONL+IS thickening observed in the SI and NI sectors in NPDR may represent early photoreceptor stress or compensatory swelling. Notably, both SwinUNETR and VM-UNet were able to replicate the general pattern of effect sizes seen in the ground-truth data, suggesting their suitability for detecting biologically meaningful trends despite minor segmentation discrepancies.
IV-B Clinical Insights
Our findings reveal significant associations between retinal layer thickness and visual acuity within NPDR and PDR patients. In the NPDR group, ground-truth analysis revealed that thickening of the RNFL (SI), OPL (SO), ONL+IS (SO), and RPE (IO) layers was significantly associated with worse vision. These findings likely reflect early microvascular and inflammatory changes, such as localized edema progressing into irreversible neurodegeneration. VM-UNet demonstrated strong concordance with the ground-truth trends, accurately capturing the association between OPL thickening in the SO sector and reduced vision, and further identifying plausible correlations in the RNFL, INL, and fluid in clinically relevant locations. In contrast, SwinUNETR did not identify any statistically significant relationships. In the PDR group, worse vision was significantly associated with thickening in the EZ (NO) and OS layers (SO, IO, TO), as well as thinning of the INL (II). These changes may indicate cumulative photoreceptor damage and fluid accumulation in the advanced DR stage. Both segmentation models identified the association between OS thickening and vision loss in the SO and IO sectors, with additional findings in the EZ and RPE layers. The results demonstrate the potential of the models to predict major vision changes via retinal structural variations.
Despite identifying biologically plausible trends, the models demonstrated limited predictive power in establishing statistically robust associations between retinal structural changes and visual acuity. While some associations reached nominal significance, none remained significant after the FDR correction. It demonstrates the inherent challenges with limited sample sizes, high-dimensional retinal imaging data, and complex model fitting procedures for generalized linear models with multiple covariates. The lack of statistical significance does not necessarily imply an absence of true effects but rather serves as a foundation for targeted hypothesis-driven studies in larger cohorts with increased statistical power.
V Limitation and Future Work
@@
First, manual labelling imperfections impact both model performance and thickness measurements. Retinal layer segmentation is inherently challenging due to subtle boundary variations and overlapping structures. In cases where excessive fluid penetrates the layer boundaries, some portions of the layer become invisible or physically diminished. More clinical expertise is needed to segment the extreme instances properly. HRF segmentation suffers from inconsistencies in ground-truth annotations, as small, widely distributed foci are challenging to delineate manually. Interestingly, in some cases, automated models provided more precise segmentations than the ground-truth labels. For example, for the first B-scan in Figure 5(a), the predictions of SegFormer and SwinUNETR have better RNFL segmentation than the ground truth with a smoother and more precise layer boundary. For the first B-scan in Figure 5(c), no fluid is manually annotated, but the segmentation models, except for SegFormer, predict potential intra-retinal fluid across the OPL and ONL+IS regions. Additionally, using more pre-processing and post-processing techniques may help improve the performance, such as the pixel-wise relative positional map as an extra input and random forest classifier as a label refiner(Ma et al., 2021).
@@
Third, this study’s cross-sectional nature limits its ability to track disease progression over time. Longitudinal studies would provide better insights into how retinal layer thickness evolves in DR. For example, several studies report RNFL/GCL thinning during the progression of DR, which has become one of the most important preclinical biomarkers for DR severity evaluation(Bhaskaran et al., 2023)(Oshitari et al., 2009)(Park et al., 2011). Additionally, while the sample size is sufficient to detect significant differences, it may limit the generalizability of the findings. A larger dataset encompassing a broader range of DR severities and treatment histories could provide more robust conclusions. In our study, no significant RNFL thickness difference is found between the NPDR and PDR groups. Fig 7 shows explicit GCL+IPL thinning of PDR in terms of the mean and interquartile range, but only the NI sector exhibits marginal significance($p=0.058$). Expanding the cohort to include more diverse patient populations may help improve the applicability of the findings across different clinical settings.
@@
Fourth, the lack of a detailed NPDR severity grading system may limit the ability to distinguish early, intermediate, and severe NPDR stages. Different NPDR severities likely exhibit distinct retinal layer changes, and a more granular classification system could enhance the understanding of DR progression. Future studies should explore integrating severity-based stratification to assess how thickness variations differ across NPDR subtypes. To our best knowledge, most DR grading datasets with public access focus on fundus color images like Messidor [^3] https://www.adcis.net/en/third-party/messidor/ and DRTiD [^4] https://github.com/FDU-VTS/DRTiD. Additional efforts are needed to investigate the OCT image associated with DR severity levels, which are precisely determined using corresponding fundus images.
@@
Lastly, integrating multi-modal imaging techniques such as OCT angiography (OCTA) could provide additional insights into the vascular changes associated with DR. For example, Alam et al. discovered the difference in vascular complexity features between NPDR and PDR patients(Alam et al., 2021). Multiple OCT parameters are significantly correlated with DR severity(Laotaweerungsawat et al., 2020). Combining structural OCT findings with functional vascular imaging may improve disease characterization and facilitate more targeted therapeutic interventions.
VI Conclusion